# A Comprehensive Survey on Underwater Image Enhancement Based on Deep Learning

### v1

@@ Abstract
Underwater image enhancement (UIE) is a challenging research task in the field of computer vision. Although hundreds of UIE algorithms have been proposed, a comprehensive and systematic review is still lacking. To promote future research, we summarize the UIE task from multiple perspectives. First, the physical models, data construction processes, evaluation metrics, and loss functions are introduced. Second, according to the contributions brought by different literatures, recent proposed algorithms are discussed and classified from six perspectives, namely network architecture, learning strategy, learning stage, assistance task, domain perspective and disentanglement fusion, respectively. Third, considering the inconsistencies in experimental settings in different literatures, a comprehensive and fair comparison does not yet exist. To this end, we quantitatively and qualitatively evaluate state-of-the-art algorithms on multiple benchmark datasets. Finally, issues worthy of further research in the UIE task are raised. A collection of useful materials is available at https://github.com/YuZhao1999/UIE.
@@ I Introduction
Underwater imaging is an important task in the field of computer vision [6, 123]. High-quality underwater images are necessary for underwater resource exploration, film shooting, personal entertainment, etc. However, due to the absorption and scattering effects of light in underwater scenes, the quality of underwater images may be degraded to varying degrees [17]. As shown in Fig. 1-(a), as the depth of water increases, the red, orange, yellow and green light components disappear in sequence. Meanwhile, as shown in Fig. 1-(b), diverse distortions may occur in underwater environments. Common types of distortion include
@@
  • Hazy, noisy and blurry effect: Light may be significantly attenuated in water, which usually caused by suspended particles or muddy water. The image obtained in such environment may appear hazy, noisy or blurry [124].
@@
Aiming at improving the quality of degraded underwater images collected in complex underwater environments, various underwater images enhancement (UIE) methods have been proposed. Existing algorithms can be divided into non-deep learning-based UIE (NDL-UIE) and deep learning-based UIE (DL-UIE). Various prior assumptions, physical models and non-data-driven classical image processing methods are widely utilized by NDL-UIE [114, 24, 115, 159, 175]. However, due to the complexity of the underwater environment, the strategy adopted by NDL-UIE may be inaccurate in certain scenarios. Problems include (i) perfect physical modeling of underwater scenarios does not exist, (ii) inherent errors exist in the estimation of physical parameters, (iii) specific prior assumptions may not applicable in every scene, and (iv) scene-specific representations not held by classical image processing methods [144, 7, 109, 58].
Considering the impressive performance of data-driven algorithms in the field of computer vision and image processing, learning-based solutions have received more attention from researchers. Aiming at further improving the performance of the UIE task, various DL-UIE algorithms have been proposed and verified [141, 53, 15, 30, 99, 122, 28, 83, 74, 105, 47]. Focusing on the problems faced by the UIE task, various efforts are expended by researchers. To promote future research, we summarize the UIE task from multiple perspectives include
  • Solutions have been proposed.
  • The results achieved by existing solutions.
@@
Solutions have been proposed. On the one hand, we summarize the physical models, data generation methods, and evaluation metrics which constructed to serve as the basis for research. On the other hand, numerous network structures and training strategies are classified and discussed. In order to know which methods have been widely explored, we need to understand the similarities and differences between them. Within our perspective, existing research can be subjectively viewed as being carried out from six different aspects, namely network architecture, learning strategy, learning stage, assistance task, domain perspective and disentanglement &\&& fusion. To facilitate future research that can easily build on existing work. According to the main contributions of different algorithms, we divide the existing work into 6 first-level categories. Then, for each first-level category, we give the corresponding second-level category. The taxonomy of algorithms is shown in Table I.
The results achieved by existing solutions. In existing papers, state-of-the-art (SOTA) algorithms have been verified on benchmark datasets. Meanwhile, their performance have been compared with other algorithms. However, the experimental settings adopted in different papers may be inconsistent. A comprehensive and fair comparative experiment for the UIE task has not yet been conducted. Here, we provide a fair comparison by unifying various experimental settings with the aim of obtaining answers on the following two questions.
@@ I-A The difference between this survey and others
The [157, 144, 109, 58, 128] mainly introduce the traditional UIE algorithms, while the latest DL-UIE algorithms are rarely mentioned. [7] classifies DL methods from the perspective of network architecture. But the latest progress in the research of the UIE task such as Fourier operation [131], contrastive learning [86], and rank learning [37] are not mentioned. The perspective of [170]’s research is the difference between hardware-based and software-based algorithms, so a comprehensive discussion of DL-based algorithms is not the goal of [170]. In this survey, we provide a comprehensive discussion of recent DL-based advancements.
@@ I-B A guide for reading this survey
The terminology used for the task of improving underwater image quality has certain differences in existing literatures, which includes underwater image enhancement, underwater image restoration and underwater image dehazing. Since certain similarities exist between the task we studied and the image dehazing task, such as similar imaging principles [11], some literature uses “underwater image dehazing” as a definition. This definition has been used less frequently in recent literature [7]. Furthermore, “enhancement” and “restoration” are often used in different literatures. There is no significant difference between these two terms in the task of improving underwater image quality. Therefore, in this survey, we use “enhancement” uniformly to define the task we discuss.
@@ II-A Problem Definition and Research Scope
where $\mathcal{L}(\hat{y},y)$ denote the loss function for obtaining the optimal parameters $\hat{\theta}$. $\mathcal{L}(\hat{y},y)$ may be supervised or unsupervised loss functions with any form.
@@ II-C Physical Models
where $\beta^{B}$ and $\beta^{D}$ represent the backscatter and direct transmission attenuation coefficients [7]. The $v_{B}$ and $v_{D}$ are the dependent coefficients for $\beta^{B}$ and $\beta^{D}$. The atmospheric scattering model and revised underwater image formation model can be used for synthesize of data and design of algorithms. Since an in-depth discussion of the imaging mechanism may be beyond the scope of this paper, more details on the physical model can be found in [2, 4, 3].
@@ II-D Datasets for Model Training and Performance Evaluation
  • Voted by SOTA UIE algorithms. 12 algorithms are used by UIEB [74] to generate diverse reference images. The best enhancement effect for each scene is then voted on by 50 volunteers as the final reference image. Meanwhile, 18 UIE algorithms and 20 volunteers are involved in the voting process of LSUI’s [102] reference images. SAUD [60] proposes a subjectively annotated benchmark that contains both real-world raw underwater images and available ranking scores under 10 different UIE algorithms.
@@
  • Collected by the undersea image capturing system. By setting up a multi-view imaging system under seawater, RUIE [85] constructs an underwater benchmark under natural light. The scenes in RUIE show a natural marine ecosystem containing various sea life, such as fish, sea urchins, sea cucumbers, scallops, etc.
@@ II-E Evaluation Metrics
The way to evaluate the performance of UIE models is a challenging topic under exploration. Currently, full-reference and no-reference evaluation metrics are the most widely used metrics. In addition, human subjective evaluation, downstream task evaluation and model efficiency are also adopted by various literatures. Details are as follows.
@@
  • No-reference metrics. For the evaluation of real-world enhancement performance without reference images, no-reference metrics are needed. The Underwater Image Quality Measures (UIQM) [101] and Underwater Color Image Quality Evaluation (UCIQE) [146] are widely adopted as no-reference metrics in literatures. Efforts [164], [60] are still being made to develop more reliable metrics for no-reference assessment.
  • Human subjective evaluation. Human vision-friendly enhanced underwater images are one of the purposes of UIE models. On the one hand, UIE-related papers [141, 53, 15, 30] usually illustrate the superiority of their proposed methods in color correction, detail restoration, and edge sharpening from the perspective of quantitative analysis. Such an evaluation is usually performed by the authors of the paper themselves. On the other hand, multiple people with or without image processing experience are involved in the evaluation of some works [74, 37]. They can give subjective ratings to different algorithms. The best enhancement result may be selected [74], or the enhancement results obtained by different algorithms may be ranked [37].
@@ II-F Loss Functions
  • L1 loss [34, 169], Smooth L1 loss [153, 166] and L2 loss [20, 111] provide the pixel-level constraint, which is used as fidelity loss.
@@
  • The attenuation of underwater images may be accompanied by the loss of edge information. So, the edge loss [51, 38, 97] calculated by the gradient operator is exploited.
@@ III UIE Methods
According to main contributions of each paper, we divide UIE algorithms into six categories, namely (i) Section III-A Network Architecture, (ii) Section III-B Learning Strategy, (iii) Section III-C Learning Stage, (iv) Section III-D Assistance Task, (v) Section III-E Domain Perspective and (vi) Section III-F Disentanglement &\&& Fusion, respectively. Details are as
@@ III-A1 Convolution Operation
Convolution is one of the most commonly used operations in UIE models. Network block constructed methods by convolution operation include naive convolution (UWCNN [72], UWNet [99]), residual connection (UResNet [83]), residual-dense connection (DUIR [25], FloodNet [32]), and multi-scale fusion (LAFFNet [143]). The computational consumption of UIE models based on convolutional architecture is usually less.
@@ III-A2 Attention Mechanism
In the UIE task, the attention mechanism is mainly used to help the model focus on different spatial locations or weight different channel information. By analyzing the wavelength-driven contextual size relationship of the underwater scene, WaveNet [111] assigns operations with different receptive fields to different color channels. Then, attention operations along spatial and channel dimensions are used to enhance convolutional paths with different receptive fields. Based on the selective kernel mechanism, a nonlinear strategy is proposed by ADMNNet [141] to change receptive field sizes reasonably by adopting soft attention. Meanwhile, a module with the ability to dynamically aggregate channel information is designed by ADMNNet. Aiming at calibrating the detailed information of the distorted input, MFEF [169] constructs a pixel-weighted channel attention calculation flow.
@@ III-A3 Transformer Module
The Transformer architecture [120, 90] has been widely used in natural language and computer vision research. UWAGA [48] points out that an automatic selection mechanism for grouping the input channels is beneficial for mining relationships between channels. Based on this view, an adaptive group attention operation embedded in Swin-Transformer [90] is designed by UWAGA. As shown in Fig. 2-(a), under the principle of Swin-Transformer, the image (or feature map) can be split into window-based patches. Image Transformers pretrained on ImageNet are used by PTT [10] to fine-tune on the UIE task, which shows that features with long-distance dependencies are effective for handling the underwater distortion. From both the channel and spatial perspectives, a multi-scale feature fusion transformer block and a global feature modeling transformer block are proposed by U-Trans [102]. A multi-domain query cascaded Transformer architecture is designed by Spectroformer [66], where localized transmission representation and global illumination information are considered. Although these Transformer-based architectures have proven beneficial for the UIE task, they often incur computational costs that exceed conventional convolution operations.
@@ III-A5 Wavelet Decomposition
where $I_{LL},I_{LH},I_{HL},I_{HH}$ denote the low-low, low-high, high-low and high-high sub-bands, respectively. The process of Wavelet decomposition of underwater images is visualized in Fig. 2-(c). UIE-WD [94] uses a dual-branch network to process images of different frequency sub-bands separately, in which each branch handles color distortion and detail blur problems respectively. By analyzing the limitations of a single pooling operation, wavelet pooling and unpooling layers based on Haar wavelets are proposed by EUIE [56]. PRWNet [49] uses wavelet boost learning to obtain low-frequency and high-frequency features. It proposes that high-frequency sub-bands represent texture and edge information, while low-frequency sub-bands contain color and lighting information. By using normalization and attention, the information of different sub-bands can be refined. A frequency subband-aware multi-level interactive wavelet enhancement module is designed by MWEN [127], aiming at building a feature extraction branch with more expressive capabilities.
@@ III-A6 Neural Architecture Search
The overall network architecture and the way modules are connected are important components of the UIE algorithm. Effective network design solutions require designers to have extensive experience and knowledge. Not only that, a large number of tentative experiments need to be performed repeatedly. Therefore, neural architecture search is explored by AutoEnhancer [116] for building UIE network structures with impressive performance. An encoder-decoder architecture is chosen by AutoEnhancer as the supernet. Three processes are used as the basis of AutoEnhancer, namely supernet training, subnet searching and subnet retraining. The first step is to obtain the parameter $W^{*}$
@@
where $S$ and $W$ denotes the searching space and network parameters for the supernet $N(S,W)$. The $L_{train}$ means the training objective function. Then the optimal subnets $S^{*}$ are searched by the trained supernet as
@@
where $Acc_{val}$ denotes the accuracy of the validation dataset. The finally optimal network parameters $W^{{}^{\prime}}$ is obtain by re-training the optimal subnet $M(\cdot)$ under data $X$ as
@@
There is currently little discussion on the automation of UIE network design and it deserves further exploration.

### v3

@@ Abstract
Underwater image enhancement (UIE) presents a significant challenge within computer vision research. Despite the development of numerous UIE algorithms, a thorough and systematic review is still absent. To foster future advancements, we provide a detailed overview of the UIE task from several perspectives. Firstly, we introduce the physical models, data construction processes, evaluation metrics, and loss functions. Secondly, we categorize and discuss recent algorithms based on their contributions, considering six aspects: network architecture, learning strategy, learning stage, auxiliary tasks, domain perspective, and disentanglement fusion. Thirdly, due to the varying experimental setups in the existing literature, a comprehensive and unbiased comparison is currently unavailable. To address this, we perform both quantitative and qualitative evaluations of state-of-the-art algorithms across multiple benchmark datasets. Lastly, we identify key areas for future research in UIE. A collection of resources for UIE can be found at https://github.com/YuZhao1999/UIE.
@@ I Introduction
Underwater imaging is an important task in the field of computer vision [6, 123]. High-quality underwater images are necessary for underwater resource exploration, film shooting, personal entertainment, etc. However, due to the absorption and scattering effects of light in underwater scenes, the quality of underwater images may be degraded to varying degrees [17]. As shown in Fig. 1-(a), as the depth of water increases, the red, orange, yellow and green light components disappear in sequence. Meanwhile, as shown in Fig. 1-(b), diverse distortions may occur in underwater environments. Common types of distortion are as follows.
@@
  • Hazy, noisy and blurry effect: Light may be significantly attenuated in water, which is usually caused by suspended particles or muddy water. The image obtained in such environment may appear hazy, noisy or blurry [124].
@@
Aiming at improving the quality of degraded underwater images collected in complex underwater environments, various underwater images enhancement (UIE) methods have been proposed. Existing algorithms can be divided into non-deep learning-based UIE (NDL-UIE) and deep learning-based UIE (DL-UIE). Various prior assumptions, physical models and non-data-driven classical image processing methods are widely utilized by NDL-UIE [114, 24, 115, 159, 175]. However, due to the complexity of the underwater environment, the strategy adopted by NDL-UIE may be inaccurate in certain scenarios. Problems include (i) the lack of perfect physical modeling for underwater scenarios, (ii) inherent errors in the estimation of physical parameters, (iii) the possibility that specific prior assumptions may not be applicable in every scene, and (iv) the limitation that classical image processing methods do not accommodate scene-specific representations [144, 7, 109, 58].
Considering the impressive performance of data-driven algorithms in the field of computer vision and image processing, learning-based solutions have received more attention from researchers. Aiming at further improving the performance of the UIE task, various DL-UIE algorithms have been proposed and verified [141, 53, 15, 30, 99, 122, 28, 83, 74, 105, 47]. Focusing on the problems faced by the UIE task, various efforts are expended by researchers. To promote future research, we summarize the UIE task from multiple perspectives, including
  • Proposed solutions.
  • Results achieved by existing solutions.
@@
Solutions have been proposed. On the one hand, we summarize the physical models, data generation methods, and evaluation metrics which are constructed to serve as the basis for research. On the other hand, numerous network structures and training strategies are classified and discussed. In order to know which methods have been widely explored, we need to understand the similarities and differences between them. Within our perspective, existing research can be subjectively viewed as being carried out from six different aspects, namely network architecture, learning strategy, learning stage, assistance task, domain perspective and disentanglement &\&& fusion. To facilitate future research that can easily build on existing work, we provide a comprehensive review of existing algorithms. According to the main contributions of different algorithms, we divide the existing work into 6 first-level categories. Then, for each first-level category, we give the corresponding second-level category. The taxonomy of algorithms is shown in Table I.
The results achieved by existing solutions. In existing papers, state-of-the-art (SOTA) algorithms have been verified on benchmark datasets. Meanwhile, their performance has been compared with other algorithms. However, the experimental settings adopted in different papers may be inconsistent. A comprehensive and fair comparative experiment for the UIE task has not yet been conducted. Here, we provide a fair comparison by unifying various experimental settings with the aim of obtaining answers on the following two questions.
@@ I-A The difference between this survey and others
The [157, 144, 109, 58, 128] mainly introduce the traditional UIE algorithms, while the latest DL-UIE algorithms are rarely mentioned. Anwar et al. [7] classify DL methods from the perspective of network architecture. However, the latest progress in the research of the UIE task such as Fourier operation [131], contrastive learning [86], and rank learning [37] are not mentioned. The perspective of [170] is the difference between hardware-based and software-based algorithms, so a comprehensive discussion of DL-based algorithms is not the goal of [170]. In this survey, we provide a comprehensive discussion of recent DL-based advancements.
@@ I-B A guide for reading this survey
The terminology used for the task of improving underwater image quality has certain differences in existing literatures, which includes underwater image enhancement, underwater image restoration and underwater image dehazing. Since certain similarities exist between the task we studied and the image dehazing task, such as similar imaging principles [11], some literature uses “underwater image dehazing” as a definition. This definition has been used less frequently in recent literature [7]. Furthermore, “enhancement” and “restoration” are often used in different literature. There is no significant difference between these two terms in the task of improving underwater image quality. Therefore, in this survey, we use “enhancement” uniformly to define the task we discuss.
@@ II-A Problem Definition and Research Scope
where $\mathcal{L}(\hat{y},y)$ denotes the loss function for obtaining the optimal parameters $\hat{\theta}$. $\mathcal{L}(\hat{y},y)$ may be supervised or unsupervised loss functions with any form.
@@ II-C Physical Models
where $\beta^{B}$ and $\beta^{D}$ represent the backscatter and direct transmission attenuation coefficients [7]. The $v_{B}$ and $v_{D}$ are the dependent coefficients for $\beta^{B}$ and $\beta^{D}$. The atmospheric scattering model and revised underwater image formation model can be used for synthesis of data and design of algorithms. Since an in-depth discussion of the imaging mechanism may be beyond the scope of this paper, more details on the physical model can be found in [2, 4, 3].
@@ II-D Datasets for Model Training and Performance Evaluation
  • Voted by SOTA UIE algorithms. Twelve algorithms are used by UIEB [74] to generate diverse reference images. The best enhancement effect for each scene is then selected by a vote of 50 volunteers as the final reference image. Meanwhile, 18 UIE algorithms and 20 volunteers are involved in the voting process of LSUI’s [102] reference images. SAUD [60] proposes a subjectively annotated benchmark that contains both real-world raw underwater images and available ranking scores under 10 different UIE algorithms.
@@
  • Collected by the undersea image capturing system. By setting up a multi-view imaging system under seawater, RUIE [85] constructs an underwater benchmark under natural light. The scenes in RUIE show a natural marine ecosystem containing various sea life, such as fish, sea urchins, sea cucumbers and scallops.
@@ II-E Evaluation Metrics
The way to evaluate the performance of UIE models is a challenging topic under exploration. Currently, full-reference and no-reference evaluation metrics are the most widely used metrics. In addition, human subjective evaluation, downstream task evaluation and model efficiency are also adopted by various literature. Details are as follows.
@@
  • No-reference metrics. For the evaluation of real-world enhancement performance without reference images, no-reference metrics are needed. The Underwater Image Quality Measures (UIQM) [101] and Underwater Color Image Quality Evaluation (UCIQE) [146] are widely adopted as no-reference metrics in literature. Efforts [164], [60] are still being made to develop more reliable metrics for no-reference assessment.
  • Human subjective evaluation. Human vision-friendly enhanced underwater images are one of the purposes of UIE models. On the one hand, UIE-related papers [141, 53, 15, 30] usually illustrate the superiority of their proposed methods in color correction, detail restoration, and edge sharpening from the perspective of quantitative analysis. Such an evaluation is usually performed by the authors of the paper themselves. On the other hand, multiple people with or without image processing experience are involved in the evaluation of [74, 37]. They can give subjective ratings to different algorithms. The best enhancement result may be selected [74], or the enhancement results obtained by different algorithms may be ranked [37].
@@ II-F Loss Functions
  • $\mathcal{L}_{1}$ loss [34, 169], Smooth $\mathcal{L}_{1}$ loss [153, 166] and $\mathcal{L}_{2}$ loss [20, 111] provide the pixel-level constraint, which is used as fidelity loss.
@@
  • The attenuation of underwater images may be accompanied by the loss of edge information. Therefore, the edge loss [51, 38, 97] calculated by the gradient operator is exploited.
@@ III UIE Methods
According to main contributions of each paper, we divide UIE algorithms into six categories, namely (i) Section III-A Network Architecture, (ii) Section III-B Learning Strategy, (iii) Section III-C Learning Stage, (iv) Section III-D Assistance Task, (v) Section III-E Domain Perspective and (vi) Section III-F Disentanglement &\&& Fusion, respectively. Details are as follows.
@@ III-A1 Convolution Operation
Convolution is one of the most commonly used operations in UIE models. Network blocks constructed by convolution operation include naive convolution (UWCNN [72], UWNet [99]), residual connection (UResNet [83]), residual-dense connection (DUIR [25], FloodNet [32]), and multi-scale fusion (LAFFNet [143]). The computational consumption of UIE models based on convolutional architecture is usually less.
@@ III-A2 Attention Mechanism
In the UIE task, the attention mechanism is mainly used to help the model focus on different spatial locations or channel information. By analyzing the wavelength-driven contextual size relationship of the underwater scene, WaveNet [111] assigns operations with different receptive fields to different color channels. Then, attention operations along spatial and channel dimensions are used to enhance convolutional paths with different receptive fields. Based on the selective kernel mechanism, a nonlinear strategy is proposed by ADMNNet [141] to change receptive field sizes reasonably by adopting soft attention. Meanwhile, a module with the ability to dynamically aggregate channel information is designed by ADMNNet. Aiming at calibrating the detailed information of the distorted input, MFEF [169] constructs a pixel-weighted channel attention calculation flow.
@@ III-A3 Transformer Module
The Transformer architecture [120, 90] has been widely used in natural language processing and computer vision research. UWAGA [48] points out that an automatic selection mechanism for grouping the input channels is beneficial for mining relationships between channels. Based on this view, an adaptive group attention operation embedded in Swin-Transformer [90] is designed by UWAGA. As shown in Fig. 2-(a), under the principle of Swin-Transformer, the image (or feature map) can be split into window-based patches. Image Transformers pretrained on ImageNet are used by PTT [10] to fine-tune on the UIE task, which shows that features with long-distance dependencies are effective for handling the underwater distortion. From both the channel and spatial perspectives, a multi-scale feature fusion transformer block and a global feature modeling transformer block are proposed by U-Trans [102]. A multi-domain query cascaded Transformer architecture is designed by Spectroformer [66], where localized transmission representation and global illumination information are considered. Although these Transformer-based architectures have proven beneficial for the UIE task, they often incur computational costs that exceed conventional convolution operations.
@@ III-A5 Wavelet Decomposition
where $I_{LL},I_{LH},I_{HL},I_{HH}$ denote the low-low, low-high, high-low and high-high sub-bands, respectively. The process of Wavelet decomposition of underwater images is visualized in Fig. 2-(c). UIE-WD [94] uses a dual-branch network to process images of different frequency sub-bands separately, in which each branch handles color distortion and detail blur problems respectively. By analyzing the limitations of a single pooling operation, wavelet pooling and unpooling layers based on Haar wavelets are proposed by EUIE [56]. PRWNet [49] uses wavelet boost learning to obtain low-frequency and high-frequency features. PRWNet proposes that high-frequency sub-bands represent texture and edge information, while low-frequency sub-bands contain color and lighting information. By using normalization and attention, the information of different sub-bands can be refined. A frequency subband-aware multi-level interactive wavelet enhancement module is designed by MWEN [127], aiming at building a feature extraction branch with more expressive capabilities.
@@ III-A6 Neural Architecture Search
The overall network architecture and the way that modules are connected are important components of the UIE algorithm. Effective network design solutions require designers to have extensive experience and knowledge. Not only that, a large number of tentative experiments need to be performed repeatedly. Therefore, neural architecture search (NAS) is explored by AutoEnhancer [116] for building UIE network structures with impressive performance. An encoder-decoder architecture is chosen by AutoEnhancer as the supernet. Three processes are used as the basis of AutoEnhancer, namely supernet training, subnet searching and subnet retraining. The first step is to obtain the parameter $W^{*}$
@@
where $S$ and $W$ denote the searching space and network parameters for the supernet $N(S,W)$, respectively. The $L_{train}$ means the training objective function. Then the optimal subnets $S^{*}$ are searched by the trained supernet as
@@
where $Acc_{val}$ denotes the accuracy of the validation dataset. The final optimal network parameters $W^{{}^{\prime}}$ is obtained by re-training the optimal subnet $M(\cdot)$ under data $X$ as
@@
There is currently little discussion on the automation of UIE network design, and it deserves further exploration.